how is 2pi equal to 360 degree and difference between radian and degree

Answer by stanbon(75887)

how is 2pi equal to 360 degree and difference between radian and degree
-------------------
Radian is a way of measuring an angle.
---
Definition: One Radian is the measure of a central subtending
an arc of the circle that is as long as the radius of the circle.
-----
How many radians are there in a circle?
The circumference of a circle = 2(pi)r
The # of radians in a circle is (2(pi)r)/r = 2(pi).
So 360 degrees = 2(pi) radians.
--------------------------
Note: One Radian is exactly 360/(2(pi)) = 180/pi degrees.
One Radian is approximately 57.3 degrees.
==============================================
Cheers,
Stan H.
===================